Fresh Foods is considering the purchase of a new packaging system. The system costs $218,071. The company plans to borrow three-quarters (3/4) of the purchase price from a bank at 9% per year compounded annually. The loan will be repaid using equal, annual payments over a 6-year period. The payments will be made at the end of each year for the life of the loan, with the first payment occurring at the end of year 1. The system is expected to last 20 years and have a salvage value of $19,521 at the end of its life.

Over the 20-year period, the company expects to pay $935 per year for maintenance. In addition, the system will require an overhaul at the end of year 8 which will cost $12,322. The system will save $4,920 per year because of efficiencies. The company uses a MARR of 4% to evaluate investments. What is the equivalent uniform annual worth (EUAW) of this system?
Answer #1

The cost of the system is mentioned as $218071 of which 75% will be financed with the bank loan. The remaining source has not been mentioned and it has been taken as initial cost.
System Cost = 218071
Loan Amount = 218071 * 0.75
= 163553.25

Remaining Cost
218071 - 163553.25 = 54517.75

Annual Payment of the Loan
=PMT(9%,6,-163553.25,,0)
= 36459.26

Now we will have to calculate PV of the loan payment, cost, salvage value and benefits with 4% as a discounting factor

PV of loan payment
=PV(4%,6,-36459.26)
= 191124.43

PV of maintenance cost
=PV(4%,20,-935)
= 12706.96

PV of benefits
=PV(4%,20,-4920)
= 66864.41

PV of overhaul cost at the end of 8th year
12322 / ( 1.04 ^ 8 ) = 7696.28

PV of salvage value
19521 / ( 1.04 ^ 20 ) = 8909.13

NPV = ( 8909.13 + 66864.41 ) - ( 54517.75 + 191124.43 + 12706.96 + 7696.28 )
= -190271.8854

PV annuity factor for 4% and 20 years
= 13.5903

Equivalent Annual Worth
= ( -190271.8854 ) / 13.5903
= -14000.54